# A Practitioner’s Guide to Kolmogorov–Arnold Networks

### v3

@@ 1 Introduction
Since then, numerous advances have broadened the scope of MLP-based PINNs. Fractional operators were addressed by fPINNs [2], while uncertainty quantification for forward and inverse problems was introduced in [3]. Extensions such as VPINNs [4], XPINNs for domain decomposition [5], and multi-fidelity PINNs [6] improved accuracy and scalability. The framework also inspired convergence analyses [7], NTK-based diagnostics [8], and practical toolkits like DeepXDE [9]. More recent contributions include multi-stage training with near machine-precision accuracy [10], adaptive residual weighting [11], and distributed solvers for extreme-scale PDEs [12, 13, 14, 15]. Collectively, these developments establish MLP–PINNs as a mature reference standard in scientific machine learning, many of which, as we show in Section 4.1, have directly inspired KAN development.
Despite their flexibility, MLPs face well-documented limitations. Fixed activation functions restrict adaptability [16]. Network behavior can be difficult to interpret [17] and to attribute causally [18]. Achieving high accuracy often entails large parameter counts, which can hinder efficient updates [19, 20]. Generalization robustness can degrade in challenging regimes [269, 21]. Optimization itself can be fragile or stiff, depending on the task and scaling [22]. MLPs also exhibit spectral bias: a tendency to learn low frequencies faster than high frequencies [23, 24]. The effect slows convergence and reduces accuracy for oscillatory or sharp-gradient solutions [25]. This motivates investigations into alternatives with explicit basis control.
KANs. Kolmogorov–Arnold Networks (KANs) in their modern multilayer form were introduced by Liu et al. [26] as a structured alternative to standard MLPs, in which fixed pointwise activations are replaced by learnable univariate functions defined on network edges and composed across layers. This design is motivated by the Kolmogorov representation theorem, but departs from its classical construction by learning the inner functions directly from data.
@@
Contributions. To the best of our knowledge, this review is:
  • the first to systematically connect KAN architectures with the Kolmogorov superposition theorem;
  • the first to comprehensively bridge KANs with classical kernel methods;
  • the first to compile and compare reported accuracy, convergence behavior, and computational cost between KANs and MLPs across tasks;
  • the first to provide a comprehensive and unified overview of basis functions used in KANs;
  • the first to collect and organize a wide range of open-source KAN implementations.
  • the most complete to date provide a structured and integrative perspective on KANs by unifying theoretical foundations, architectural designs, optimization strategies, and applications into a coherent roadmap for both understanding and practice.
@@
Table 1: Representative performance trends comparing KAN-based models with MLPs. Results are aggregated from the literature and reflect general outcomes. “Slower training” refers to higher per-iteration cost, though total wall-clock time may differ due to fewer epochs to convergence. All rows compare KANs to plain MLP/PINN baselines without advanced features.
@@
[72] PIKAN $\approx$ PINNs Slower training B-spline
@@
…/KAN_PointNet_CFD Jacobi-based network for CFD predictions. [72]
@@ 2 Bridging KANs and KST
The construction of $\psi$ involves three parameters with clearly distinct roles. The dimension parameter $n$ determines how many input variables are being folded into one and therefore directly controls how complex the encoding must be. The base parameter $\gamma$ sets the scale separation of the digit-based representation. The refinement level $k$ specifies how many recursive steps are applied. Increasing $k$ refines the same basic geometric pattern, while changing $n$ modifies the structure of that pattern itself.
@@ 3 Bridging KANs and Kernel Methods
    The distinction is therefore explicit. In the classical case, the mapping from data to coefficients is linear and fully described by the design (kernel) matrix $\Psi$. In the deep KAN, each coefficient $C_{k}$ is a nonlinear function of parameters from multiple layers. Terms such as ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}^{2}$ arise from composition and appear in several coefficients simultaneously, meaning that a single parameter influences multiple polynomial orders. Consequently, increasing depth not only raises the effective polynomial degree, but also couples the coefficients nonlinearly, making their identification a fundamentally different problem from solving a linear kernel system, even in one-dimensional settings.
@@ 4.1 How KANs Extend the MLP Landscape
Vision and Representation Learning. In computer vision, Convolutional KANs embed spline-based activations directly within convolutional kernels, producing more expressive yet lightweight CNNs [172] [^5] https://github.com/AntonioTepsich/Convolutional-KANs. Residual KAN modules integrate into ResNet backbones to improve gradient flow and generalization [177] [^6] https://github.com/withray/residualKAN, while U-KAN extends U-Nets for image segmentation and diffusion models [179] [^7] https://github.com/CUHK-AIM-Group/U-KAN. Similarly, the KAN-Mixer adapts MLP-Mixer architectures for image classification [192] [^8] https://github.com/engichang1467/KAN-Mixer. Additional applications include remote sensing [181], hyperspectral imaging via Wav-KAN [180], medical image classification [182], and autoencoding tasks [178] [^9] https://github.com/SekiroRong/KAN-AutoEncoder.
@@ 5.1 B-spline
To match boundary and interior regularity/support, the knot vector is often extended by $k$ points on each side. This padding allows the model to progressively capture finer details [55]. As illustrated in Figure 2, the number of active basis functions grows from GkG-k to $G+k$, where $G$ is the number of internal intervals in the original, non–extended grid.
@@ 5.3 ReLU
Figure 4: (a) Step-by-step construction of the normalized ReLU-based local basis function from its constituent $\operatorname{ReLU}(e_{i}-x)$ and $\operatorname{ReLU}(x-s_{i})$ terms. (b) Complete set of normalized ReLU–KAN basis functions $R_{i}(x)$ with supports $s_{i},e_{i}] [$129].
@@ 5.5 Gaussian (RBF)
As researchers sought to improve the efficiency and performance of the original B–spline–based KANs, the Gaussian radial basis function (RBF) emerged as a powerful and computationally elegant alternative. Its smooth, localized, and infinitely differentiable nature makes it an excellent choice for representing univariate activation functions along the edges of a KAN. Across recent studies [125, 56, 148, 126, 239, 122], Gaussian–based activations have appeared in several formulations—ranging from fixed–grid hybrids to fully learnable and reflectional variants—each striking a different balance between stability, adaptivity, and efficiency.
@@
Pure RBF Layers with Learnable Centers. A more flexible RBF formulation appears in the DeepOKAN architecture of Abueidda et al. [56] [^26] https://github.com/DiabAbu/DeepOKAN, where Gaussian bases are used in both branch and trunk networks. Each activation takes the form
@@
allowing the RBFs to migrate toward regions of high variation. In contrast to some later operator-learning studies, Abueidda et al. report that RBF–KANs resolve high-frequency sinusoidal fields with markedly higher accuracy than MLPs, confirming the strong approximation capability of Gaussian bases in mechanics problems.
A complementary conclusion arises in the recent two-step DeepOKAN for bubble dynamics [149]: while RBFs remain effective in the branch network, spline bases outperform them in the trunk for extremely high-frequency, high-amplitude oscillations. Together, these works show that RBFs are highly competitive for moderate-to-high frequency regimes, but basis choice in KAN operator networks remains problem-dependent.
@@
Practical caveat: the shape parameter. Gaussian RBF performance is highly sensitive to the width parameter ($\varepsilon$), which governs both smoothness and conditioning [206, 207, 208]. Even small changes can strongly impact accuracy and numerical stability [209, 210, 211]. From a kernel–theoretic perspective, this sensitivity reflects an uncertainty principle: larger shape (width) parameters produce flatter, globally correlated kernels that improve smoothness and accuracy but worsen conditioning, whereas smaller parameters yield well–conditioned, highly localized bases at the expense of approximation power [204, 205]. While heuristic tuning and learnable widths (as in FastKAN and DeepONet–RBF variants) offer partial remedies, robust and theoretically guided selection of this parameter remains an active research problem at the intersection of kernel methods and neural approximation.
@@ 5.6 Fourier
Fourier features provide global, smooth, periodic expressivity that complements localized bases (e.g., B–splines, RBFs). Within the KAN family, Fourier–based KANs replace the univariate edge functions with harmonic expansions, retaining the KAN wiring while improving spectral resolution on tasks with strong periodic structure [140]. By contrast, Fourier networks such as FAN are not KANs: they embed cosine–sine transforms directly into each layer as a drop-in MLP replacement, targeting efficient periodicity modeling (often with reduced parameters/FLOPs) and stronger in-/out-of-domain generalization [143] [^32] https://github.com/YihongDong/FAN. We mention FAN here as a related spectral approach; our focus remains on Fourier features used inside KANs.
@@ 5.7 Wavelet
Wavelet activations in KAN layers. Wav-KAN embeds analytic wavelet functions directly into univariate edge functions. Each activation is parameterized by scale $s$ and translation $u$:
@@
Applications further highlight the versatility of wavelet bases: Pratyush et al. [218] employ a DoG wavelet in CaLMPhosKAN for enhanced phosphorylation-site detection, while Seydi et al. [180] combine CWT and DWT for hyperspectral imaging, outperforming spline-KANs and MLPs.
@@ 6.2 Adaptive Sampling & Grids
    ensuring that endpoints remain fixed ($t_{0}=a$, $t_{n}=b$). This naturally places knots in regions of high variation or non-smoothness.
@@
Grid–adaptive PIKANs. Rigas et al. [124, 267] extend this idea by adapting not only the spline grid but also the collocation points, so resolution follows the PDE residual:
@@ 6.4 Function decomposition
The intuition is that the LF model captures the dominant structure, while the HF nonlinear head corrects only the residual. This reduces overfitting and stabilizes training when HF data are scarce. MFKAN achieved lower relative $L_{2}$ errors than single-fidelity KANs under the same HF budget [195]. The same fidelity split also benefits PIKANs, where $K_{L}$ serves as a coarse physics surrogate and the HF head is trained only on residual or boundary losses, making fidelity decomposition complementary to spatial domain decomposition (Sec. 6.3).
@@ 6.5 Hybrid, Ensemble, and Data-Integrated Models
  • MLP–KAN hybrids. These architectures use MLPs for broad feature extraction and KANs for structured refinement. Examples include the MLP–KAN mixture-of-experts of He et al. [43] and the HPKM-PINN parallel fusion model of Xu et al. [60].
@@ 6.6 Sequence and Attention Hybrids
  • Attention-enhanced KANs. These models insert KAN blocks into attention layers to improve multi-scale representation. Includes KAN–MHA [58], FlashKAT (Group-Rational KAN in Transformers) [199], and Attention-KAN-PINN for battery SOH forecasting [215].
@@ 6.7 Discontinuities and Sharp Gradients
  • Physics-driven shock stabilization. Embeds stabilization mechanisms from numerical PDEs directly into the loss. Includes cPIKAN+EVM/RBA [39], Buckley–Leverett KANs [63], and EPi–cKAN for elasto-plastic transitions [64].
@@ 6.8 Optimization & Adaptive Training
Initialization and depth-stabilized training. Beyond optimizer choice and probabilistic tuning, Rigas et al. [270] [^44] https://github.com/srigas/RGA-KANs show that deep Chebyshev-based KANs (cPIKANs) fail primarily due to variance blow-up across layers. They introduce a basis-agnostic Glorot-like initialization that preserves forward/backward variance and a Residual–Gated Adaptive block that stabilizes deep gradient flow. For spline KANs, Rigas et al. [283] [^45] https://github.com/srigas/KAN_Initialization_Schemes demonstrate that Glorot-style and power-law initializations substantially improve NTK conditioning and convergence across regression and PDE benchmarks, underscoring initialization as a key design lever for scalable KAN depth.

### v4

@@ 1 Introduction
Within the modern machine learning landscape, multilayer perceptrons (MLPs) remain a foundational architecture for function approximation and data-driven modeling. Recently, Kolmogorov–Arnold Networks (KANs), introduced by Liu et al. [1, 2], have emerged as a promising alternative that rethinks the role of nonlinearities in neural networks. While MLPs rely on fixed activation functions applied at nodes, KANs instead place learnable univariate functions along network edges, a design loosely inspired by the Kolmogorov–Arnold representation theorem. This structural shift has been argued to improve interpretability and adaptability, although empirical studies suggest that its advantages over well-designed MLPs depend strongly on the problem setting and architectural choices. This emphasis on interpretability has also extended to symbolic regression and equation discovery, where KAN-based formulations aim to produce sparser, more structured, and more human-readable representations [3, 4]. To place these developments in context, we first review the role of MLPs and their extensions in scientific machine learning before introducing the KAN framework, whose conceptual origins arise from a different theoretical tradition.
@@
Since then, numerous advances have broadened the scope of MLP-based PINNs. Fractional operators were addressed by fPINNs [6], while uncertainty quantification for forward and inverse problems was introduced in [7]. Extensions such as VPINNs [8], XPINNs for domain decomposition [9], and multi-fidelity PINNs [10] improved accuracy and scalability. The framework also inspired convergence analyses [11], NTK-based diagnostics [12], and practical toolkits like DeepXDE [13]. More recent contributions include multi-stage training with near machine-precision accuracy [14], adaptive residual weighting [15], and distributed solvers for extreme-scale PDEs [16, 17, 18, 19]. Collectively, these developments establish MLP–PINNs as a mature reference standard in scientific machine learning. At the same time, KANs arise from a distinct line of thought rooted in functional representations inspired by the Kolmogorov superposition theorem, rather than as a direct modification of the MLP paradigm.
Despite their flexibility, MLPs face well-documented limitations. Fixed activation functions restrict adaptability [20]. Network behavior can be difficult to interpret [21] and to attribute causally [22]. Achieving high accuracy often entails large parameter counts, which can hinder efficient updates [23, 25, 24]. Generalization robustness can degrade in challenging regimes [277, 26]. Optimization itself can be fragile or stiff, depending on the task and scaling [27]. MLPs also exhibit spectral bias: a tendency to learn low frequencies faster than high frequencies [28, 29]. The effect slows convergence and reduces accuracy for oscillatory or sharp-gradient solutions [30]. This motivates investigations into alternatives with explicit basis control.
KANs. Kolmogorov–Arnold Networks in their modern multilayer form were introduced by Liu et al. [1] as a structured alternative to standard MLPs, in which fixed pointwise activations are replaced by learnable univariate functions defined on network edges and composed across layers. This design is motivated by the Kolmogorov representation theorem, but departs from its classical construction by learning the inner functions directly from data.
@@
Contributions. This review makes the following contributions:
  • it systematically connects KAN architectures with the Kolmogorov superposition theorem;
  • it comprehensively bridges KANs with classical kernel methods;
  • it compiles and compares reported accuracy, convergence behavior, and computational cost between KANs and MLPs across tasks;
  • it provides a comprehensive and unified overview of basis functions used in KANs;
  • it collects and organizes a wide range of open-source KAN implementations;
  • it offers a structured and integrative perspective on KANs by unifying theoretical foundations, architectural designs, optimization strategies, and applications into a coherent roadmap for both understanding and practice.
@@
Table 1: Representative performance trends comparing KAN-based models with MLPs. Results are aggregated from the literature and indicate qualitative trends. Reported accuracy and convergence metrics differ across studies (e.g., relative $L_{2}$, RMSE). “Slower training” refers to higher per-iteration cost; total time may vary with epochs. All rows compare KANs with plain MLP/PINN baselines.
@@
[75] PIKAN $\approx$ PINNs Slower training Chebyshev
@@
…/PIKAN_PointNet Chebushev for solving time-independent inverse problems. [75]
@@ 2 Bridging KANs and KST
The construction of $\psi$ involves three parameters with clearly distinct roles. The dimension parameter $n$ determines how many input variables are being folded into one and therefore directly controls how complex the encoding must be. The base parameter $\gamma>1$ denotes the scaling factor used in the digit-based representation that separates successive refinement levels. The refinement level $k\in\mathbb{N}$ indicates the number of recursive construction steps used to generate the finite approximation $\psi_{k}$ of the inner function. Increasing $k$ refines the same basic geometric pattern, while changing $n$ modifies the structure of that pattern itself.
@@ 3 Bridging KANs and Kernel Methods
    The distinction is therefore explicit. In the classical case, the mapping from data to coefficients is linear and fully described by the design (kernel) matrix $\Psi$. In the deep KAN, each coefficient $C_{k}$ is a nonlinear function of parameters from multiple layers. Terms such as $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}$} arise from composition and appear in several coefficients simultaneously, meaning that a single parameter influences multiple polynomial orders. Consequently, increasing depth not only raises the effective polynomial degree, but also couples the coefficients nonlinearly, making their identification a fundamentally different problem from solving a linear kernel system, even in one-dimensional settings.
@@ 4.1 How KANs Extend the MLP Landscape
Vision and Representation Learning. In computer vision, Convolutional KANs embed spline-based activations directly within convolutional kernels, producing more expressive yet lightweight CNNs [175] [^5] https://github.com/AntonioTepsich/Convolutional-KANs and [305]. Residual KAN modules integrate into ResNet backbones to improve gradient flow and generalization [180] [^6] https://github.com/withray/residualKAN, while U-KAN extends U-Nets for image segmentation and diffusion models [182] [^7] https://github.com/CUHK-AIM-Group/U-KAN. Similarly, the KAN-Mixer adapts MLP-Mixer architectures for image classification [195] [^8] https://github.com/engichang1467/KAN-Mixer. Additional applications include remote sensing [184], hyperspectral imaging via Wav-KAN [183], medical image classification [185], and autoencoding tasks [181] [^9] https://github.com/SekiroRong/KAN-AutoEncoder.
@@ 5.1 B-spline
Variants of this formulation also employ fractional B-splines. Aghaei and Zaky [310] introduce a physics-informed fractional KAN in which fractional B-spline bases are combined with a tensorized operational-matrix formulation to efficiently solve distributed-order fractional differential equations. These variants extend the classical spline basis to fractional operators but retain the same underlying KAN framework.
For the standard spline construction used in most KAN implementations, the knot vector is often extended by $k$ points on each side. This padding allows the model to progressively capture finer details [58]. As illustrated in Figure 2, the number of active basis functions grows from GkG-k to $G+k$, where $G$ is the number of internal intervals in the original, non–extended grid.
@@ 5.3 ReLU
Figure 4: (a) Step-by-step construction of the normalized ReLU-based local basis function from its constituent $\operatorname{ReLU}(e_{i}-x)$ and $\operatorname{ReLU}(x-s_{i})$ terms. (b) Complete set of normalized ReLU–KAN basis functions $R_{i}(x)$ with supports $s_{i},e_{i}] [$132].
@@ 5.5 Gaussian (RBF)
As researchers sought to improve the efficiency and performance of the original B–spline–based KANs, the Gaussian radial basis function (RBF) emerged as a powerful and computationally elegant alternative. Its smooth, localized, and infinitely differentiable nature makes it an excellent choice for representing univariate activation functions along the edges of a KAN.
Across recent studies [128, 59, 151, 129, 247, 125, 304, 314, 317], Gaussian–based activations have appeared in several formulations—ranging from fixed–grid hybrids to fully learnable and reflectional variants—each striking a different balance between stability, adaptivity, and efficiency.
@@
Pure RBF Layers with Learnable Parameters. A more flexible RBF formulation appears in the DeepOKAN architecture of Abueidda et al. [59] [^26] https://github.com/DiabAbu/DeepOKAN, where Gaussian bases are used in both branch and trunk networks. Each activation takes the form
@@
allowing the RBFs to migrate toward regions of high variation. Abueidda et al. show that RBF–KANs resolve high-frequency sinusoidal fields with markedly higher accuracy than MLPs in mechanics problems. A similar observation appears in the two-step DeepOKAN for bubble dynamics [152], where RBFs remain effective in the branch network while spline bases outperform them in the trunk for extremely high-frequency oscillations.
Beyond operator-learning settings, the Free-RBF-KAN model of Chiu et al. [317] introduces adaptive radial basis functions whose centers and widths evolve during training, improving efficiency and stability in general function-learning tasks.
@@
Practical caveat: the shape parameter. Gaussian RBF performance is highly sensitive to the width parameter ($\varepsilon$), which governs both smoothness and conditioning [209, 210, 211, 212, 213, 214, 215, 216]. Even small changes can strongly impact accuracy and numerical stability [217, 218, 219]. From a kernel–theoretic perspective, this sensitivity reflects an uncertainty principle: larger shape (width) parameters produce flatter, globally correlated kernels that improve smoothness and accuracy but worsen conditioning, whereas smaller parameters yield well–conditioned, highly localized bases at the expense of approximation power [207, 208]. While heuristic tuning and learnable widths (as in FastKAN and DeepONet–RBF variants) offer partial remedies, robust and theoretically guided selection of this parameter remains an active research problem at the intersection of kernel methods and neural approximation.
@@ 5.6 Fourier
Fourier features provide global, smooth, periodic expressivity that complements localized bases (e.g., B–splines, RBFs). Within the KAN family, Fourier–based KANs replace the univariate edge functions with harmonic expansions, retaining the KAN wiring while improving spectral resolution on tasks with strong periodic structure [143, 296, 311]. By contrast, Fourier networks such as FAN are not KANs: they embed cosine–sine transforms directly into each layer as a drop-in MLP replacement, targeting efficient periodicity modeling (often with reduced parameters/FLOPs) and stronger in-/out-of-domain generalization [146] [^32] https://github.com/YihongDong/FAN. We mention FAN here as a related spectral approach; our focus remains on Fourier features used inside KANs.
@@ 5.7 Wavelet
Wavelet activations in KAN layers. Wav-KAN embeds analytic wavelet functions directly into univariate edge functions. Each activation is parameterized by scale $s$ and translation $u$ [148, 297, 312]:
@@
Applications further highlight the versatility of wavelet bases: Pratyush et al. [226] employ a DoG wavelet in CaLMPhosKAN for enhanced phosphorylation-site detection, while Seydi et al. [183] combine CWT and DWT for hyperspectral imaging, outperforming spline-KANs and MLPs. Wavelet–KAN blocks have also been integrated with attention mechanisms for dense prediction tasks such as medical image segmentation [313].
@@ 6.2 Adaptive Sampling & Grids
    ensuring that endpoints remain fixed ($t_{0}=a$, $t_{n}=b$). This naturally places knots in regions of high variation or non-smoothness. Related analyses of free-knot spline behavior in KANs have been carried out by Zheng et al. [316] [^40] https://github.com/IcurasLW/FR-KAN, who derive bounds on the number of spline knots and propose a Free-Knots KAN that improves training stability while reducing the number of trainable parameters.
@@
Grid–adaptive PIKANs. Rigas et al. [127, 275] extend this idea by adapting not only the spline grid but also the collocation points, so resolution follows the PDE residual. More recently, a dynamic grid-adaptation framework was proposed that treats knot allocation as a density-estimation problem governed by importance density functions, allowing curvature or training metrics to determine grid resolution during learning [315].
@@ 6.4 Function decomposition
The intuition is that the LF model captures the dominant structure, while the HF nonlinear head corrects only the residual. This reduces overfitting and stabilizes training when HF data are scarce. MFKAN achieved lower relative $L_{2}$ errors than single-fidelity KANs under the same HF budget [198].
The same fidelity split also benefits PIKANs, where $K_{L}$ serves as a coarse physics surrogate and the HF head is trained only on residual or boundary losses, making fidelity decomposition complementary to spatial domain decomposition (Sec. 6.3). Related coupled PIKAN strategies sequentially train subnetworks representing different subsystems to stabilize optimization [302], while multi-branch physics-informed KANs decompose interacting physical subsystems into parallel KAN branches with physics constraints [308].
The same fidelity split also benefits PIKANs, where $K_{L}$ serves as a coarse physics surrogate and the HF head is trained only on residual or boundary losses, making fidelity decomposition complementary to spatial domain decomposition (Sec. 6.3).
@@ 6.5 Hybrid, Ensemble, and Data-Integrated Models
  • Function-combination KANs. FC-KAN combines multiple basis families (e.g., B-splines, wavelets, and radial basis functions) through element-wise operations such as sums, products, and concatenations, enabling richer functional representations within a single KAN layer [299] [^42] https://github.com/hoangthangta/FC_KAN.
  • MLP–KAN hybrids. These architectures use MLPs for broad feature extraction and KANs for structured refinement. Examples include the MLP–KAN mixture-of-experts of He et al. [46], the HPKM-PINN parallel fusion model of Xu et al. [63], and two-stage medical analysis pipelines that couple foundation-model segmentation with KAN-based classification [306].
@@
  • Symbolic-regression hybrids. Howard et al. [3] propose SINDy-KANs, which combine KANs with SINDy-style sparse regression to identify parsimonious symbolic representations at the level of the learned activation functions. This hybrid design targets equation discovery and dynamical-system identification, showing that KANs can be coupled not only with neural modules, but also with sparse scientific model-discovery frameworks.
@@ 6.6 Sequence and Attention Hybrids
  • Attention-enhanced KANs. These models insert KAN blocks into attention layers to improve multi-scale representation. Includes KAN–MHA [61], FlashKAT (Group-Rational KAN in Transformers) [202], Attention-KAN-PINN for battery SOH forecasting [223], and wavelet-attention hybrids such as DWSen [312].
@@ 6.7 Discontinuities and Sharp Gradients
  • Physics-driven shock stabilization. Includes cPIKAN+EVM/RBA [42], Buckley–Leverett KANs [66], EPi–cKAN for elasto-plastic transitions [67],and physics-informed KAN solvers for the Vlasov–Poisson system in cosmological structure formation [301].
@@ 6.8 Optimization & Adaptive Training
Initialization and depth-stabilized training. Beyond optimizer choice and probabilistic tuning, Rigas et al. [278] [^46] https://github.com/srigas/RGA-KANs show that deep Chebyshev-based KANs (cPIKANs) fail primarily due to variance blow-up across layers. They introduce a basis-agnostic Glorot-like initialization that preserves forward/backward variance and a Residual–Gated Adaptive block that stabilizes deep gradient flow.
For spline KANs, Rigas et al. [291] [^47] https://github.com/srigas/KAN_Initialization_Schemes demonstrate that Glorot-style and power-law initializations substantially improve NTK conditioning and convergence across regression and PDE benchmarks, underscoring initialization as a key design lever for scalable KAN depth. Related efficiency improvements can also be achieved through feature enrichment strategies such as FEKAN [309], which augment the input representation before the KAN mapping to accelerate convergence without modifying the underlying architecture.